OpenStudy (curry):

a roller coaster at an amusement park is at rest of top of a 30m hill (point A). the car starts to roll down the hill and reaches (point b), which is 10 m above the ground, and then rolls up the track to point C, which is 20 m above the ground. - A student assumes no energy is lost, and solves for how fast is the car moving at point C using energy arguements. What answer does he get?

OpenStudy (anonymous):

use potential and kinetic energy equations. Set them equal to each other Then you can cancel out mass

OpenStudy (curry):

I got 10rad2 but the answer is 20 m/s. I got 20m/x at point B not point C. For point C, I got 10rad2 = rad 200

OpenStudy (anonymous):

you only have to find the energy at point C. the kinetic energy at point C will be the same as the potential energy in the difference of height from point A to point C. does that make any sense?

OpenStudy (anonymous):

I don't have the actual answers.

OpenStudy (anonymous):

yet

OpenStudy (curry):

it does, but the answer still shouldn't be 20m/s

OpenStudy (anonymous):

and I don't think it is 20m/s at point B

OpenStudy (anonymous):

it's not 20 m/s mgh=mv^2 gh=v^2

OpenStudy (curry):

KE = 0.5mv^2

OpenStudy (anonymous):

oh. yes. thank you.

OpenStudy (anonymous):

2gh=v^2

OpenStudy (curry):

in which case, I'm right? and the book is wrong?

OpenStudy (anonymous):

yes

OpenStudy (curry):

kk thanks!! :D